# Masked Thought: Simply Masking Partial Reasoning Steps Can Improve Mathematical Reasoning Learning of Language Models

### v1

@@ 3.2 Main Results
MaskedLM $seq_{i}=\begin{cases}\text{[mask]}&\text{with probability }m,\\ w\sim\text{Uniform}(\text{Vocabulary})&\text{with probability }r,\\ \end{cases}$ m=1,r=0 46.7
@@ 3.3 Comparison with Other Regularization Techniques
There are inquiries into whether alternative approaches can yield comparable outcomes. Thus, we delve into different techniques for introducing noise and empirically evaluate their effectiveness. We explore the effects of adding noises to different positions, as well as the noise type introduced by NEFTune, Dropout, Scheduled Sampling, and MaskedLM (our primary implementation). Detailed descriptions of these methods are provided in the Appendix K.
Results. Our experiments yield the following insights: 1. In terms of the placement of noise, we find that masking partial reasoning steps in the target yields the best performance when compared to introducing noise across all tokens or specifically within the question. 2. We observe that introducing more huge noise to remove the original semantic is advantageous. Specifically, setting $\tau=100$ for Scheduled Sampling and $a=1000$ for NEFTune enhances the model’s performance. In contrast, the strategy of replacing synonyms, with $\tau=1,2$ in Scheduled Sampling, does not surpass the effectiveness of random replacement.
When the model is trained with MaskedLM using $r=1$, the model encounters random tokens unrelated to the current context, leading it to disregard the semantics of these tokens, akin to the role of [mask]. In the experiments in Tab. 3, $r=1$ yields slightly higher results than $m=1$. While $r=1$ demonstrates higher levels of randomness in comparison to $m=1$, it has the potential to better mitigate overfitting. However, on average, the distinction between $r=1$ and $m=1$ is marginal for larger datasets. With the exception of using $r=1$ for GSM8K-7K, we default to utilizing $m=1$ for other datasets in Tab. 1. We also explore the impact of adding attention masking at the masked positions within each transformer layer, beyond just the input masking. We find that its effectiveness is on par with using input masking alone.
Additionally, we observe that even when solely computing losses on non-noisy tokens in MFT, the performance still surpasses that of standard SFT (45.0, +3.4). Nonetheless, it exhibits lower performance compared to the computation of losses across all tokens (47.1, +5.5). Predicting the next tokens for masked positions poses a greater challenge than for unmasked ones, as they lack adjacent unmasked tokens from which to gather local information. This increased complexity ultimately improves the performance.
@@ 3.4 Investigating the Impact of MFT on Language Modeling
Case Study. To supplement the statistical results shown in Figure 1, we further conduct a case study to demonstrate the impact of token dependencies introduced by MFT. In Figure 4, we vary a digit within the context between 00 and $9$ to see whether it influences the predictions of subsequent steps. If the model consistently predicts $16-3$, this indicates that the model’s prediction is not influenced by the digit change, but rather relies on the information provided in the question and prior steps, with the word “three” highlighted in orange. Conversely, if the model’s prediction shifts to $16-n$, it implies a dependency on immediate, adjacent tokens, suggesting a reliance on distance shortcuts. The frequency of $16-3$ predictions, as summarized in Table 4, reveals that MFT primarily leverages the token within the question, in contrast to the SFT approach, which tends to adopt a more short-sighted view.
Table 4: Results of model predictions after altering. When alter n to 0-9, we provide the proportion of the model predicting 16-3 or 16-n. A larger number represents less sensitive to the nearby premise and stronger dependence on questions.
@@
Conversely, we also modify the premise within the question as shown in Fig. 5, where a prediction of $16-n$ indicates that, when faced with conflicting premises, the model prioritizes the question. The results of altering in question demonstrate that MFT excels in interpreting the questions, while SFT tends to ignore changes within the question.
Table 5: We conduct tests on GSM-IC (Grade-School Math with Irrelevant Context) Shi et al. (2023) to compare the ability to cope with the distractors in the questions. This result illustrates that MFT can more effectively eliminate the distractors, utilizing more useful premises from the questions.
@@
Dealing with Distractors. Some previous studies have also indicated that adding some disturbance to questions, such as inserting a irrelevant context Shi et al. (2023) or changing the order of premises within the question Chen et al. (2024), can decrease the reasoning ability of large language models. We test on the GSM-IC Shi et al. (2023) dataset which inserts some irrelevant distractor sentences into the original questions in GSM8K. From the results in Tab. 5, it can be seen that although MFT and RFT are roughly comparable on the original dataset, MFT shows a greater improvement on the GSM-IC dataset. This suggests that there are differences in the mechanisms by which these two methods improve performance and indicates that MFT can effectively locate useful premises within the disturbed questions, even though MFT does not mask the questions during training.
Table 6: Using GPT-4 to split the answer into steps and judge the correctness of middle steps in 200 samples. False Positive means the predicted answer is correct but the correct steps is less than total steps.
@@
MFT ($m$=0,$r$=1) 3 75 0.540 3.2
False Positive Analysis. Despite the premises in these experiments being artificially disturbed, we would like to know if MFT might generate false positives due to the reduced use of neighboring features. False positives in this context refer to instances where intermediate steps are incorrect, yet the final result is accurately predicted. This analysis is conducted using GPT-4, which achieves an accuracy of 92% on the specified dataset. The findings, as detailed in Table 6, indicate that MFT does not result in a significant rise in the occurrence of false positives. The model’s autoregressive prediction process not only improves the accuracy of the final answer but also preserves the correctness of the intermediate steps. We provide the prompt of GPT-4 and the cases of false positives in Appendix F.
Figure 6: We conduct experiments across two datasets to investigate the impacts of datasets on dependency. Left: Llama-GSM8K. Right: Llama-Alpaca. Training on Alpaca does not enhance the dependency, unlike GSM8K.
Impact of Datasets on Learning Dependency. We investigate the impact of different datasets on the effectiveness of the MFT technique, specifically, which types of data are best suited for MFT in Fig. 6. Our experiments focus on the Alpaca dataset Taori et al. (2023), a general instruction set with minimal mathematical reasoning content. The results reveal significant differences between the Alpaca and GSM8K datasets. Training with MFT using the Alpaca dataset does not alter dependency relationships, verifying our hypothesis. MFT enhances the model’s ability to learn long-distance dependencies within the data. The absence of such dependencies in data renders MFT less effective. Additionally, we assess whether MFT could detrimentally influence the model’s general conversational capabilities. According to the results presented in Appendix H, the scores of Alpaca-eval Li et al. (2023b) using both MFT and SFT were comparable.
Method OSME ISE & SUE CE Total Errors
SFT 15.6% 78.0% 6.0% 173
RFT-100 11.9% 77.4% 10.6% 151
MFT 15.2% 74.8% 9.9% 151
Error Analysis. We also perform a quantitative error analysis to understand the impact of MFT on the final results by analyzing the types of errors. We adopt the following error types: Calculator Error (CE), One Step Missing Error (OSME), Semantic Understanding Error (SUE) and Incoherent Steps Error (ISE). CE and OSME are minor errors that can be fixed with one-step while SUE and ISE are errors of misunderstanding the question or previous steps which need substantial fixing as defined by Wei et al. (2022). See more details of these error types in Appendix J. We use GPT-4 as a proxy for the judgement. Tab. 7 demonstrates that MFT has fewer SUE and ISE errors. This may be due to the benefits obtained by enhancing the model’s question dependence. Additionally, the model experiences an increase in OSME and CE compared to SFT. The increased ratios might be minor errors that arise after solving major errors in question understanding.
@@ 4 Conclusion
This method is effective for reasoning data, while demonstrating no significant effect on a general instruction tuning dataset which lacks of multi-steps reasoning. We have only provided a qualitative explanation so far, without proposing a method that includes quantitative metrics to determine which types of data are best suited for this approach.
@@
Acknowledgments
This work was supported by Alibaba Group through Alibaba Innovative Research Program.
@@ Appendix A Related work
A.1 Mathematical Reasoning of LLM.
To further enhance the capabilities of pre-trained Large Language ModelsTouvron et al. (2023); Jiang et al. (2023); Achiam et al. (2023); Azerbayev et al. (2023) on reasoning tasks, the primary strategies involve fine-tuning and various training-free post-processing techniques. The post-processing techniques include approaches like promptingWei et al. (2022); Fu et al. (2022); Kojima et al. (2022); Wang et al. (2023a), ensemblingWang et al. (2022); Li et al. (2022), and rerankingUesato et al. (2022); Cobbe et al. (2021). A unique technique specific to LLMs is to leverage the model’s reasoning capabilities through prompting. This entails crafting zero-shot promptsKojima et al. (2022); Wang et al. (2023a) or designing a small set of demonstrations to facilitate in-context learningWei et al. (2022); Fu et al. (2022). Additionally, there are methods focused on designing intricate pipelines utilizing various rules or algorithms for enhanced reasoning. These methods include searchYao et al. (2023); Besta et al. (2023), iterative feedback collectionYao et al. (2022), and employing the Program-of-Thought approach for generating code solutionsChen et al. (2022).
In terms of fine-tuning, whether through Supervised Fine Tuning (SFT) or Reinforcement Learning (RL), the primary aim is to gather high-quality supervisory signals for model improvement. For SFT, the approach generally involves leveraging larger models or manual efforts to collect high-quality data An et al. (2023); Yu et al. (2023); Yue et al. (2023); Luo et al. (2023). MetaMathYu et al. (2023); Gou et al. (2023) augments its dataset by rewriting mathematical queries and provide answers with the help of an LLM. It also leverages backward reasoning to enhance the model’s capabilities. MAmmoTHYue et al. (2023) compiles diverse mathematical datasets for training, enriching them with Chains of Thought (CoT) and Programs of Thought (PoT) rationales for reasoning. RL-based methods strive to gather superior training signals through self-sampling and reward supervision.  Lightman et al. (2023) proposes to use process supervision to give more precise signals than outcome supervision. Some strategies adopt the process supervision into the online RL training Wang et al. (2023b); Xi et al. (2024); Luo et al. (2023). RFTYuan et al. (2023) is a simplified offline method of standard online RL. It enhances its training dataset with accurate reasoning paths by generating and collecting them via a SFT model. This approach is further extended by ReST Singh et al. (2023) through an iterative process.
@@
A.2 Data Augmentation for NLP.
In the area of Natural Language Processing (NLP), an extensive range of data augmentation techniques has been introduced for diverse applications, such as translation, summarization, and dialogue Chen et al. (2023b); Fadaee et al. (2017); Xie et al. (2016); Lample et al. (2017); Guo et al. (2020a); Xie et al. (2022); Wang et al. (2018). The noise-based strategies commonly involve the integration of random continuous noise Cheng et al. (2018); Jain et al. (2023); Kedzie and McKeown (2019), token dropout Xie et al. (2016); Gal and Ghahramani (2015), token swapping Wei and Zou (2019), and more advanced mixup strategies Guo et al. (2020b); Xie et al. (2022). Recently, in the context of regularization for LLMs, Jain et al. (2023) has demonstrated that the application of continuous noise across all tokens during the instruction tuning can significantly improve the length and detail of the generated responses. However, this technique does not improve the reasoning ability. We adopt the discrete token noise through token dropout Xie et al. (2016); Gal and Ghahramani (2015) to reasoning tasks, finding that it works well, which may complement the dependency-rich structure characteristic of reasoning data.
@@ B.2 Reinforcement Learning (RL) Fine-tuning
After SFT, to further improve the model, RL is applied to align the model to the specific goal. The model automatically samples responses, and a reward model scores these generated responses to distinguish their quality. For mathematical reasoning, it has been verified that rewards based on intermediate processes yield better results compared to rewards based solely on final results Wang et al. (2023b); Lightman et al. (2023).
@@ B.4 Discussion
Question: Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May? Answer1: 1. Natalia sold 48/2 = 24 clips in May. 2. Natalia sold 48+24 = 72 clips altogether in April and May. 3. The answer is 72 Answer2: 1. Natalia sold 48/2 = 24 clips in May. 2. She sold 48+24 = 72 clips altogether in April and May. 3. The answer is 72 Answer3: 1. Natalia sold 48/2=24 clips in May. 2. The total number of clips sold by Natalia in April and May is 48+24=72 3. The answer is 72 Answer4: 1. The number of clips sold in May is 1/2 * 48 = 24. 2. Altogether, Natalia sold 48 + 24 = 72 clips in April and May. 3. The answer is 72
@@
Sample Efficiency. For online RL, it is necessary to sample diverse and better reasoning paths to achieve better results than using a static dataset. However, after SFT for initialization, the sampled paths will be very similar to the ground truth. If the sampling is sufficient and the reward model is strong enough to find high-quality reasoning paths that may have better generality and conciseness, RL can achieve better results. This can be the reason why RFT which using much more samples performs similarly to ours, as RFT only samples for multiple paths at once and only uses outcome supervision.
@@
(a) A reasoning procedure represented by a graph of thought.
(b) Representing the graph by a chain of thought, the model is expected to learn the the reasoning relationship denoted by the blue solid lines.
(c) The solution space forms a fully connected graph within the Transformer architecture.
(d) Language models may overfit on the local information while neglect other premises.
(e) We randomly prune the fully-connected graph to compel the model to leverage potentially distant premises.
Figure 8: The model acquires the ability to engage in step-by-step reasoning by leveraging prior premises through data-driven learning. We add regularization to language modeling, compelling the model to grasp the deductive relationship inherent between premises and conclusions.

### v2

@@ 3.2 Main Results
MaskedLM $seq_{i}=\begin{cases}\text{[mask]}&\text{prob }m,\\ w\sim\mathcal{U}(V)&\text{prob }r,\\ \end{cases}$ m=1,r=0 46.7
@@ 3.3 Comparison with Other Regularization Techniques
There are inquiries into whether alternative approaches can yield comparable outcomes. Thus, we delve into different techniques for introducing noise and empirically evaluate their effectiveness. We explore the effects of adding noises to different positions, as well as the noise type introduced by NEFTune, Dropout, Scheduled Sampling, and MaskedLM (our primary implementation). The variables of compared methods are introduced in Tab. 4. Detailed descriptions of these methods are provided in the Appendix K.
Results. Our experiments yield the following insights: 1. In terms of the placement of noise, we find that masking partial reasoning steps in the target yields the best performance when compared to introducing noise across all tokens or specifically within the question. 2. We observe that introducing more huge noise to remove the original semantic is advantageous. Specifically, setting the temperature $\tau=100$ for Scheduled Sampling and the noise scale $a=1000$ for NEFTune enhances the model’s performance. In contrast, the strategy of replacing synonyms with $\tau=1,2$ in Scheduled Sampling does not surpass the effectiveness of random replacement.
When the model is trained with MaskedLM using $r=1$, i.e., replacing with a random token, the model encounters tokens unrelated to the current context, leading it to disregard the semantics of these tokens, akin to the role of [mask]. In the experiments in Tab. 5, $r=1$ yields slightly higher results than $m=1$. While $r=1$ demonstrates higher levels of randomness in comparison to $m=1$, it has the potential to better mitigate overfitting. However, on average, the distinction between $r=1$ and $m=1$ is marginal for larger datasets. With the exception of using $r=1$ for GSM8K, we default to utilizing $m=1$ for other datasets in Tab. 1. We also explore the impact of adding attention masking at the masked positions within each transformer layer beyond just the input masking. We find that its effectiveness is on par with that of using input masking alone.
Another hypothesis is that the model could benefit from predicting multiple tokens ahead when doing MFT, so we conduct an additional comparison. We observe that even when solely computing losses on non-noisy tokens in MFT where no multiple-token prediction occurs, the performance still surpasses that of standard SFT (45.0, +3.4). Nonetheless, it exhibits lower performance compared to the computation of losses across all tokens (47.1, +5.5). Predicting the next tokens for masked positions poses a greater complexity than for unmasked ones, as they lack adjacent unmasked tokens from which to gather local information. We believe this increased complexity further improves the performance.
@@ 3.4 Investigating the Impact of MFT on Language Modeling
Case Study. We conduct case studies on token dependencies to show that MFT helps the model to use correct premises even when misleading errors occur in the chain of thoughts. we observe that the dependency changes are consistent across different samples and present one as an example . In Fig. 4, we vary a digit within the context between 00 and $9$ to see whether it influences the predictions of subsequent steps. If the model consistently predicts $16-3$, this indicates that the model’s prediction is not influenced by the digit change but rather relies on the information provided in the question and prior steps, with the word “three” highlighted in orange. Conversely, if the model’s prediction shifts to $16-n$, it implies a dependency on immediate, adjacent tokens, suggesting a reliance on distance shortcuts. The frequency of $16-3$ predictions, as summarized in Tab. 6, reveals that MFT establishes longer dependencies to locate correct premises rather than relying only on recent intermediate results in the chain of thoughts.
Table 6: Results of model predictions after altering. When altering n to 0-9, we provide the proportion of the model predicting 16-3 or 16-n. A larger number indicates being less sensitive to the nearby premise and stronger dependence on questions.
@@
Conversely, we also modify the premise within the question as shown in Fig. 5, where a prediction of $16-n$ indicates that, when faced with conflicting premises, the model prioritizes the question. The results of altering the question demonstrate that MFT excels in interpreting the questions, while SFT tends to ignore changes within the question.
Table 7: We conduct tests on GSM-IC Shi et al. (2023) to compare the ability to cope with the distractors in the questions. This result illustrates that MFT can more effectively eliminate the distractors, utilizing useful premises from the questions.
@@
Dealing with Distractors. Some previous studies have also indicated that adding some disturbance to questions, such as inserting an irrelevant context Shi et al. (2023) or changing the order of premises within the question Chen et al. (2024), can decrease the reasoning ability of large language models. We test on the GSM-IC Shi et al. (2023) dataset, which inserts some irrelevant distractor sentences into the original questions in GSM8K. The results in Tab. 7 show that although MFT and RFT are roughly comparable on the original dataset, MFT shows a greater improvement on the GSM-IC dataset. This suggests that there are differences in the mechanisms by which the two methods improve performance and indicates that MFT can effectively locate useful premises within the disturbed questions, even though MFT does not mask the questions during training.
Table 8: Using GPT-4 to split the answer into steps and judge the correctness of middle steps in 200 samples. False Positive means the predicted answer is correct, but the correct steps are less than the total steps.
@@
MFT 3 75 0.540 3.2
False Positive Analysis. Despite the premises in these experiments being artificially disturbed, we would like to know if MFT might generate false positives due to the reduced use of neighboring features. False positives in this context refer to instances where intermediate steps are incorrect, yet the final result is accurately predicted. This analysis is conducted using GPT-4666To help GPT-4 judge the correctness of solutions, we include the ground-truth reasoning steps in the prompts. We provide the prompt of GPT-4 and the cases of detected false positives in Appendix F., which achieves an accuracy of 92% on the specified dataset. The findings, as detailed in Tab. 8, indicate that MFT does not result in a significant rise in the occurrence of false positives. MFT not only improves the accuracy of the final answer but also preserves the correctness of the intermediate steps.
4 Related work
Mathematical Reasoning of LLM. To improve the mathematical reasoning ability through supervised fine-tuning, the approach generally involves leveraging larger models or manual efforts to collect high-quality data An et al. (2023); Yu et al. (2023a); Yue et al. (2023); Luo et al. (2023); Tao et al. (2024); Gou et al. (2023). MetaMath Yu et al. (2023a) augments its dataset by rewriting mathematical queries and providing answers with the help of an LLM. MAmmoTH Yue et al. (2023) compiles diverse mathematical datasets for training, enriching them with Chains of Thought (CoT) and Programs of Thought (PoT) rationales for reasoning, resulting in the MathInstruct Dataset. RFT Yuan et al. (2023) is a simplified offline method of standard online RL. It enhances its training dataset with accurate reasoning paths by generating and collecting them via an SFT model.
Data Augmentation for NLP. Noise-based strategies are widely used in the field of data augmentation for NLP. Common strategies include the integration of random continuous noise Cheng et al. (2018); Jain et al. (2023); Kedzie and McKeown (2019), token dropout Xie et al. (2016); Gal and Ghahramani (2015), token swapping Wei and Zou (2019), and more advanced mixup strategies Guo et al. (2020a); Xie et al. (2022). We adopt the discrete token noise through token dropout Xie et al. (2016); Gal and Ghahramani (2015) to reasoning tasks, finding that it works well, which may complement the dependency-rich structure characteristic of reasoning data.
We provide a more thorough discussion of related works in Appendix A.
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (NSFC Grant No. 62122089), Beijing Outstanding Young Scientist Program NO. BJJWZYJH012019100020098, Intelligent Social Governance Platform, Major Innovation & Planning Interdisciplinary Platform for the "Double-First Class" Initiative, Renmin University of China, the Fundamental Research Funds for the Central Universities, the Research Funds of Renmin University of China, and Alibaba Group through Alibaba Innovative Research Program.
@@ Acknowledgments
This method is effective for reasoning data while demonstrating no significant effect on a general instruction tuning dataset which lacks multi-step reasoning. We have only provided a qualitative explanation so far without proposing a method that includes quantitative metrics to determine which types of data are best suited for this approach.
@@ Appendix A Related Work
A.1 Mathematical Reasoning of LLM
Pre-trained Large Language Models have demonstrated impressive reasoning capabilities and have shown corresponding improvements on various downstream tasks Kojima et al. (2022); Huang et al. (2022); Yang et al. (2024); Wang et al. (2023c); Liu et al. (2024). To further enhance the capabilities of pre-trained Large Language Models Touvron et al. (2023); Jiang et al. (2023); Achiam et al. (2023); Azerbayev et al. (2023) on reasoning tasks, the primary strategies involve fine-tuning and various training-free post-processing techniques. The post-processing techniques include approaches like prompting  Wei et al. (2022); Fu et al. (2022); Kojima et al. (2022); Wang et al. (2023a); Gao et al. (2024), ensembling  Wang et al. (2022b); Li et al. (2022), and reranking  Uesato et al. (2022); Cobbe et al. (2021). A unique technique specific to LLMs is to leverage the model’s reasoning capabilities through prompting. This entails crafting zero-shot prompts  Kojima et al. (2022); Wang et al. (2023a) or designing a small set of demonstrations to facilitate in-context learning  Wei et al. (2022); Fu et al. (2022). Additionally, there are methods focused on designing intricate pipelines utilizing various rules or algorithms for enhanced reasoning. These methods include search  Yao et al. (2023); Besta et al. (2023), iterative feedback collection  Yao et al. (2022), and employing the Program-of-Thought approach for generating code solutions Chen et al. (2022b).
In terms of fine-tuning, whether through Supervised Fine Tuning (SFT) or Reinforcement Learning (RL), the primary aim is to gather high-quality supervisory signals for model improvement. For SFT, the approach generally involves leveraging larger models or manual efforts to collect high-quality data An et al. (2023); Yu et al. (2023a); Yue et al. (2023); Luo et al. (2023); Tao et al. (2024); Gou et al. (2023). MetaMath  Yu et al. (2023a) augments its dataset by rewriting mathematical queries and providing answers with the help of an LLM. It also leverages backward reasoning to enhance the model’s capabilities. MAmmoTH Yue et al. (2023) compiles diverse mathematical datasets for training, enriching them with Chains of Thought (CoT) and Programs of Thought (PoT) rationales for reasoning. RL-based methods strive to gather superior training signals through self-sampling and reward supervision.  Lightman et al. (2023) proposes to use process supervision to give more precise signals than outcome supervision. Some strategies adopt the process supervision into the online RL training Wang et al. (2023b); Xi et al. (2024); Luo et al. (2023); Yu et al. (2023b). RFT Yuan et al. (2023) is a simplified offline method of standard online RL. It enhances its training dataset with accurate reasoning paths by generating and collecting them via an SFT model. This approach is further extended by ReST Singh et al. (2023) through an iterative process.
@@ A.1 Mathematical Reasoning of LLM
A.2 Data Augmentation for NLP
In the area of Natural Language Processing (NLP), an extensive range of data augmentation techniques has been introduced for diverse applications, such as translation, summarization, and dialogue Chen et al. (2023b); Fadaee et al. (2017); Xie et al. (2016); Lample et al. (2017); Guo et al. (2020b); Xie et al. (2022); Wang et al. (2018); Lin et al. (2022); Chen et al. (2022a). The noise-based strategies commonly involve the integration of random continuous noise Cheng et al. (2018); Jain et al. (2023); Kedzie and McKeown (2019), token dropout Xie et al. (2016); Gal and Ghahramani (2015), token swapping Wei and Zou (2019), and more advanced mixup strategies Guo et al. (2020a); Xie et al. (2022). Recently, in the context of regularization for LLMs, Jain et al. (2023) has demonstrated that the application of continuous noise across all tokens during the instruction tuning can significantly improve the length and detail of the generated responses. However, this technique does not improve the reasoning ability. We adopt the discrete token noise through token dropout Xie et al. (2016); Gal and Ghahramani (2015) to reasoning tasks, finding that it works well, which may complement the dependency-rich structure characteristic of reasoning data.
@@ Appendix B Discussion of the Relation to RL
Since MFT and RFT exhibit similar performance, and RL is also considered a highly effective method for reasoning tasks, we discuss the relationship between MFT and RL and try to explain the reasons for the similar performance of MFT and RFT.
@@ B.2 Reinforcement Learning (RL) Fine-tuning
After SFT, to further improve the model, RL is applied to align the model to the specific goal Wang et al. (2023b, 2022a, 2021). The model automatically samples responses, and a reward model scores these generated responses to distinguish their quality. For mathematical reasoning, it has been verified that rewards based on intermediate processes yield better results compared to rewards based solely on final results Wang et al. (2023b); Lightman et al. (2023).
@@ B.4 Discussion
Example: Sampled Correct Reasoning Paths of a SFT Model Question: Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May? Answer1: 1. Natalia sold 48/2 = 24 clips in May. 2. Natalia sold 48+24 = 72 clips altogether in April and May. 3. The answer is 72 Answer2: 1. Natalia sold 48/2 = 24 clips in May. 2. She sold 48+24 = 72 clips altogether in April and May. 3. The answer is 72 Answer3: 1. Natalia sold 48/2=24 clips in May. 2. The total number of clips sold by Natalia in April and May is 48+24=72 3. The answer is 72 Answer4: 1. The number of clips sold in May is 1/2 * 48 = 24. 2. Altogether, Natalia sold 48 + 24 = 72 clips in April and May. 3. The answer is 72
@@
Sample Efficiency. For online RL, it is necessary to sample diverse and better reasoning paths to achieve better results than using a static dataset. However, after SFT for initialization, the sampled paths of the trained query will be very similar to the ground truth. If the sampling is sufficient and the reward model is strong enough to find high-quality reasoning paths that may have better generality and conciseness, RL can achieve better results. This can be the reason why RFT, which uses much more samples, performs similarly to ours, as RFT only samples for multiple paths at once and only uses outcome supervision. On the contrary, online RL can gain more benefits Wang et al. (2023b).